OpenStudy (anonymous):

A used book store started selling used CDs and videos. In the first eek the store sold 40 used CDs and videos at 4$ per CD and 6$ for a Video. The sales for both CDs and VIdeos totaled 180$. 1:Write a system of equations to represent the situation.

OpenStudy (anonymous):

Lets call the number of CDs 'c' and the number of Videos 'v'. From the question we know that: The total number of items sold is 40: c+v=40 We also know that the total money made is $180 and the price each item sells for: 4c+6v=180 So our system of equations is: c + v = 40 4c+ 6v= 180